For the directed weighted graph given below find shortest distances and shortest paths from A to all other vertices. Use the Dijkstra algorithm. Show the status of the array of distances after each iteration of the while loop.

2-1 C ) ? H e- 90? 2 2 22 (4-21?121? G
Answer #1
A B C D E F G H I
0 \infty \infty \infty \infty \infty \infty \infty \infty
0 1 \infty 2 \infty \infty \infty \infty \infty
0 1 \infty 2 3 \infty \infty \infty \infty
0 1 \infty 2 3 \infty 4 \infty \infty
0 1 5 2 3 5 4 5 5
0 1 5 2 3 5 4 5 5
0 1 5 2 3 5 4 5 5
0 1 5 2 3 5 4 5 5
0 1 5 2 3 5 4 5 5
0 1 5 2 3 5 4 5 5